OpenStudy (anonymous):

Use the Distributive Property and simplify the expression. 4 + 2(3x – 1) + 9 A. 6x + 11 B. 6x + 15 C. 5x + 14 D. 3x + 14

OpenStudy (anonymous):

il give medal

OpenStudy (anonymous):

plz

OpenStudy (adajiamcneal):

A, do you need the work

OpenStudy (anonymous):

yes

OpenStudy (adajiamcneal):

ok hold on and excuse the ugly hand writing xD

OpenStudy (anonymous):

plz can you help wth a few more

OpenStudy (adajiamcneal):



OpenStudy (adajiamcneal):

and ill try

OpenStudy (anonymous):

Which is 1024 written as a power of a number? A. 2564 B. 5122 C. 45 D. 54

OpenStudy (adajiamcneal):

do any of the answers have exponents?

OpenStudy (anonymous):

c and i wrote it wrong :p

OpenStudy (adajiamcneal):

what?

OpenStudy (anonymous):

4^5

OpenStudy (adajiamcneal):

wait so you already have the answer?

OpenStudy (anonymous):

ya for that one

OpenStudy (adajiamcneal):

ok then whats the next question

OpenStudy (anonymous):

What is the value of 54 – 4a2 + 3b3 when a = −2 and b = 4?

OpenStudy (adajiamcneal):

plug in a and b and then use pemdas

OpenStudy (adajiamcneal):

is 2 the power of 4a?

OpenStudy (adajiamcneal):

\[54 - 4 ^{-2} + 3\times4 ^{3}\]

OpenStudy (anonymous):

help

OpenStudy (adajiamcneal):

haha im solving it rn

OpenStudy (anonymous):

lol

OpenStudy (adajiamcneal):

what are your options cuz I think im wrong

OpenStudy (anonymous):

really

OpenStudy (adajiamcneal):

yeah I got 833 so whats your options for that equation

OpenStudy (adajiamcneal):

or I googled it and got 106 id go with google if I were you haha

OpenStudy (triciaal):

@adajiamcneal you have -2 the original has 2

OpenStudy (adajiamcneal):

ohhhhh ok xD thank you haha

OpenStudy (anonymous):

106 ?

OpenStudy (adajiamcneal):

I would say yes because I keep getting higher answers so just take googles answer haha

OpenStudy (adajiamcneal):

next question?

OpenStudy (triciaal):

54 - 16 + 192 = 230

OpenStudy (anonymous):

what is the value of b ?b^3 1 over 216

OpenStudy (adajiamcneal):

that is 6

OpenStudy (anonymous):

Which is equal to 3^3 • 3^5?

OpenStudy (triciaal):

b^-3 = 1/b^3 = 1/216 b^3 = 216 take the cube root of each side (b^3)^1/3 = (216)^1/3 yes b = 6

OpenStudy (anonymous):

3^8 ?

OpenStudy (adajiamcneal):

yeah I was actually solving it pero that is right too haha

OpenStudy (anonymous):

12^5=12^?

OpenStudy (adajiamcneal):

uhh obviously if the first equation is 12 to the 5th power and it has an equal sign then it would be the same thing haha aka 12^5 is the answer

OpenStudy (anonymous):

12^5^3=12^?

OpenStudy (anonymous):

@adajiamcneal

OpenStudy (adajiamcneal):

ohhhh 12 ^8

OpenStudy (anonymous):

@adajiamcneal

OpenStudy (adajiamcneal):

give me uno momento xD

OpenStudy (adajiamcneal):

idk on that one either sorry :/

OpenStudy (anonymous):

What is the standard form of the number shown in the calculator display? 5.93e-5

OpenStudy (adajiamcneal):

593000

OpenStudy (yanasidlinskiy):

The last one is....... -'s go to the left. So go 10 decimals to the left. and on the other one, go 6 decimals to the left:)

OpenStudy (adajiamcneal):

you already asked me that one I couldn't solve it haha

OpenStudy (anonymous):

new question

OpenStudy (anonymous):

@adajiamcneal

OpenStudy (adajiamcneal):

160

OpenStudy (anonymous):

Combine the like terms to simplify the expression: 5 + 7 + 15xy2 + 9y3 + 10x –13xy2 – 3x – 8y3 @adajiamcneal

OpenStudy (adajiamcneal):

12+ 2xy^2 + 1y^3+13x

OpenStudy (anonymous):

A. y6 + 2x2y2 + 10x + 12 B. y3 –2xy2 + 10x + 15 C. y3 + 2xy2 + 7x + 12 D. y3 + 2xy2 + 10x + 1

OpenStudy (anonymous):

these are the answers

OpenStudy (adajiamcneal):

y^3 + 2xy^2 + 7x + 12

OpenStudy (adajiamcneal):

c

OpenStudy (anonymous):

thx il tellu how much i got

OpenStudy (adajiamcneal):

ok haha wish you the best :)

OpenStudy (anonymous):

@adajiamcneal

OpenStudy (adajiamcneal):

6 ^2

OpenStudy (anonymous):

A. 616 B. 64 C. 6−6 D. 6−10

OpenStudy (adajiamcneal):

D

OpenStudy (anonymous):

a.6^16 b.6^4 c.6^-6 d.6^10

OpenStudy (adajiamcneal):

I still believe its d because it has to be ten

OpenStudy (anonymous):

66%

OpenStudy (adajiamcneal):

dammmn im sorry /.\ at least its not an f

OpenStudy (anonymous):

lol your right :/

OpenStudy (adajiamcneal):

next time ill just google all the answers pero goodnight :)

OpenStudy (anonymous):

night

OpenStudy (triciaal):

12^5^3=12^? this is not 12^8 it would be 12^15

OpenStudy (triciaal):

6^-8/6^-2 = 6^(-8--2) = 6^(-8 + 2) = 6^-6 = C